The population of a town is increasing at a rate of 6.2% per year. The city council believes they will have to add another elementary school when the population reaches 100 000. If there are currently 76 000 people living in the town, how long do they have before the new school will be needed?
I can't remember what equation to sub it into, but the textbook answer is 4.6 years.

Answer by stanbon(75887)

The population of a town is increasing at a rate of 6.2% per year. The city council believes they will have to add another elementary school when the population reaches 100 000. If there are currently 76 000 people living in the town, how long do they have before the new school will be needed?
I can't remember what equation to sub it into, but the textbook answer is 4.6 years.
------
y = ab^x
-------
100,000 = 76,000*(1.062)^x
------
(1.062)^x = 25/19
-----
x = log(25/19)/log(1.062)
-----
x = 4.5622
=============
Cheers,
Stan H.
=============